# Analyzing Fairness of Computer Vision and Natural Language Processing Models

### v2

@@ Abstract
Machine learning (ML) algorithms play a crucial role in decision-making across diverse fields such as healthcare, finance, education, and law enforcement. Despite their widespread adoption, these systems raise ethical and social concerns due to potential biases and fairness issues. This study focuses on evaluating and improving the fairness of Computer Vision and Natural Language Processing (NLP) models applied to unstructured datasets, emphasizing how biased predictions can reinforce existing systemic inequalities. A publicly available dataset from Kaggle was utilized to simulate a practical scenario for examining fairness in ML workflows.
To address and mitigate biases, the study employed two leading fairness libraries: Fairlearn by Microsoft, and AIF360 by IBM. These tools offer comprehensive frameworks for fairness analysis, including metrics evaluation, result visualization, and bias mitigation techniques. The research aims to measure bias levels in ML models, compare the effectiveness of these fairness libraries, and provide actionable recommendations for practitioners.
The results demonstrate that each library possesses distinct strengths and limitations in evaluating and mitigating fairness. By systematically analyzing these tools, the study contributes valuable insights to the growing field of ML fairness, offering practical guidance for integrating fairness solutions into real-world applications. This research underscores the importance of building more equitable and responsible machine learning systems.
@@ 1 Introduction
Machine learning algorithms are extensively employed across various fields, including entertainment, shopping, healthcare, finance, education, law enforcement, and critical areas such as loan approvals [1] and hiring decisions [2, 3]. They offer benefits like consistent performance and the ability to analyze numerous variables [4, 5]. However, these systems are not immune to biases, which can lead to unfair outcomes [6, 7]. Biases in machine learning are particularly concerning when decisions directly affect individuals or communities, potentially leading to discriminatory results. Ensuring that these systems operate ethically and equitably is crucial. Fair decision-making necessitates impartiality, avoiding favoritism based on inherent or acquired characteristics. Biased algorithms deviate from this principle, skewing decisions in favor of certain groups.
@@
The idea of ”fairness” in algorithmic systems is deeply tied to the sociotechnical context. Different forms of fairness-related harms have been identified:
@@
Allocation Harm: Unequal distribution of opportunities or resources, such as an algorithm disproportionately favoring men over women for job offers [8].
@@
Quality-of-Service Harm: Unequal performance across groups, e.g., facial recognition systems misclassifying Black women more frequently than White men [9], or speech recognition systems underperforming for users with speech disabilities [10].
@@
Stereotyping Harm: Reinforcing societal stereotypes, such as image searches for ”CEO” showing predominantly White men [8].
@@
Denigration Harm: Producing offensive outputs, such as misclassifying individuals as animals or chatbots generating slurs [8].
5.
Representation Harm: Over- or under-representing certain groups, such as racial bias in welfare fraud investigations or neglecting elderly populations in surveillance applications [8].
6.
Procedural Harm: Violations of social norms in decision-making practices, such as penalizing job candidates for extensive experience or failing to ensure transparency and accountability in algorithmic decisions [11].
These harms often overlap and highlight the importance of addressing fairness concerns from the development phase of machine learning systems.
@@
This study examines the fairness of machine learning models applied to structured datasets in classification tasks. Structured datasets, known for their organized format, facilitate efficient analysis and are frequently used in ML applications. However, these datasets often reflect historical biases or systemic discrimination, which can influence the fairness of machine learning models, making robust evaluation and mitigation essential.
@@
For this research, a publicly available dataset [12] from Kaggle was chosen. Kaggle datasets offer realistic scenarios for evaluating machine learning models and are well-suited for fairness studies. The dataset underwent preprocessing and was used to develop classification models, which involve predicting discrete labels based on input features. Classification tasks are particularly significant in fairness research since biased predictions can disproportionately impact specific groups.
@@
To identify and address biases, the study employed three advanced fairness libraries: Fairlearn by Microsoft, AIF360 by IBM, and the What-If Tool by Google. These libraries provide tools for assessing fairness metrics, visualizing model behavior, and applying bias mitigation techniques. Leveraging these tools, the research systematically evaluates the fairness of classification models and investigates strategies to minimize bias.
The objectives of this study are:
@@
To measure the level of bias present in machine learning models trained on the chosen dataset.
@@
3.
To offer actionable insights on implementing fairness tools in real-world machine learning workflows.
The remainder of this paper is structured as follows: Section 2 reviews related work on machine learning fairness. Section 3 details the methodology, including dataset preprocessing and model development. Section 4 discusses fairness analyses using the selected libraries and their functionalities. Section 5 presents a comparative evaluation and results of the libraries. Finally, Section 6 concludes with a summary of the findings.
2 Review of Related Work
Bias in machine learning (ML) models is an increasingly pressing concern, particularly as these models influence critical societal sectors like healthcare, hiring, and criminal justice. A growing body of research has examined the origins, manifestations, and mitigation strategies of bias, offering a solid foundation for tackling this pervasive issue.
One prominent area of study focuses on identifying and categorizing biases in ML models. Ref. [13] outlines a taxonomy of biases, classifying them into historical, representation, and measurement biases. Historical bias arises from pre-existing inequities in the data, even before applying ML methods. Representation bias occurs when certain groups are under- or over-represented in the training data, leading to skewed predictions [14]. Measurement bias results from inaccuracies in the features or labels used for training due to flawed measurement processes.
Another line of research addresses methods for detecting bias. Tools such as disparate impact analysis [15] and fairness metrics like demographic parity, equal opportunity, and disparate mistreatment [16] are widely used. In structured datasets, researchers focus on two fairness dimensions: group fairness, which ensures equitable treatment across demographic groups, and individual fairness, which emphasizes treating similar individuals similarly [17]. Ref. [18] explores the conflicts between these fairness definitions, highlighting the need for trade-offs tailored to specific contexts.
The technical challenges of mitigating bias have also been extensively studied. Pre-processing techniques, such as re-weighting data samples or modifying labels to enhance fairness, address biases at the dataset level [19]. In-processing methods, such as adversarial debiasing [20], embed fairness constraints directly into model training. Post-processing approaches modify model outputs to meet fairness criteria, including re-ranking methods proposed by [16]. However, Ref. [21] illustrates the inherent trade-offs between fairness and accuracy, demonstrating the difficulty of optimizing both simultaneously.
Bias analysis in structured datasets has received significant attention due to the prevalence of tabular data in decision-making systems. Such datasets often carry latent biases reflecting historical inequities or systemic discrimination. For instance, the COMPAS dataset, used in criminal justice, exhibits racial disparities in predictive outcomes [22]. Research has also highlighted how feature selection and preprocessing can either mitigate or exacerbate biases. Ref. [23] discusses the impact of feature correlation with sensitive attributes on fairness and proposes methods to disentangle these relationships.
Recent advances emphasize interpretability as a critical component of bias analysis. Ref. [24] introduced LIME (Local Interpretable Model-agnostic Explanations), enabling stakeholders to understand and detect biased patterns in model predictions. Similarly, Ref. [25] developed SHAP (SHapley Additive exPlanations), providing consistent and accurate feature importance values. These interpretability tools have been instrumental in identifying biases in structured datasets by allowing detailed analysis of how features contribute to unfair outcomes.
Intersectionality has also emerged as a key consideration in bias studies. Ref. [26] highlights the necessity of evaluating models across multiple demographic dimensions, revealing compounded disparities for intersectional groups, such as Black women in facial recognition systems. For structured datasets, Ref. [27] proposes fairness-enhancing strategies that account for multiple subgroups simultaneously, avoiding the limitations of single-axis fairness analyses.
The literature on bias in machine learning encompasses a broad spectrum of topics, from foundational concepts and detection methods to mitigation strategies and interpretability tools. Despite significant progress, challenges remain in applying these techniques to structured datasets, particularly in balancing fairness with competing goals like accuracy and interpretability. This review emphasizes the need for continued research into holistic, context-sensitive approaches for understanding and addressing bias in ML models.
3 Dataset and Model Details
For the computer vision model, we used the UTKFace dataset [12]. The UTKFace dataset is a collection of over 20,000 facial images labeled with demographic information, including age, gender, and ethnicity. It is widely used for machine learning tasks such as facial analysis and demographic prediction. The dataset includes images cropped and aligned to facilitate consistent analysis across models.
In terms of machine learning applications, this dataset is commonly utilized for tasks like age and gender prediction, facial recognition, and demographic analysis. Researchers often use deep learning techniques, including convolutional neural networks (CNNs), to extract features and train models on this dataset for predictive tasks.
For the Natural Language Processing (NLP) model, we used the California Independent Medical Review (IMR) Dataset [12]. The California Independent Medical Review (IMR) Dataset available on Kaggle contains healthcare-related data focusing on medical review decisions. It is primarily used for Natural Language Processing (NLP) tasks and contains text-based reviews related to insurance appeals. The dataset is valuable for modeling tasks that involve text classification, sentiment analysis, or decision-making predictions in a healthcare context.
While the dataset itself consists of text entries, researchers and developers may employ machine learning models such as Logistic Regression, Random Forest, or advanced NLP models like BERT for tasks such as understanding decision outcomes or identifying patterns in appeals based on textual data.
4 Implementation of Fairness Analyses
In recent years, the growing use of machine learning (ML) across various sectors has raised significant concerns regarding fairness and bias in classification models. Given their potential impact on decision-making in critical fields such as healthcare, finance, and criminal justice [28], ensuring fairness in these models has become crucial. Bias in ML models can arise from factors like imbalanced training data, model selection, and inherent societal biases, ultimately resulting in discriminatory outcomes against marginalized groups [29].
@@ 4 Implementation of Fairness Analyses
To address these challenges, several libraries and tools have been created to help practitioners analyze and reduce bias in their models. Notable among them are Fairlearn, AIF360, and the What-If Tool, each offering unique features for evaluating and enhancing fairness.
Fairlearn [30]: Developed by Microsoft, this toolkit assists data scientists in assessing and improving the fairness of AI models. It provides a range of metrics to evaluate fairness and algorithms to mitigate unfairness. Fairlearn emphasizes both the social and technical dimensions of fairness, helping users understand how various aspects of a model contribute to disparities among groups.
AIF360 [31]: Developed by IBM, AIF360 is a comprehensive library that offers a variety of metrics for fairness assessment and techniques for mitigating bias across the entire AI application lifecycle. It supports integration at various stages of the machine learning pipeline, promoting fairness-aware modeling. The library includes metrics for evaluating fairness across societal demographics and offers re-parameterization strategies to improve model robustness.
What-If Tool [32]: Created by Google, this interactive visualization tool enables users to explore and analyze ML models without needing to write any code. It facilitates testing model performance in hypothetical scenarios, helping users understand and explain model behavior. The tool allows users to examine how different demographics influence model predictions, providing a deeper understanding of potential bias.
These tools are vital for researchers aiming to identify and mitigate bias in ML classification models, offering the methodologies necessary to ensure equitable AI systems and informed decision-making processes.

### v3

@@ Abstract
Machine learning (ML) algorithms play a critical role in decision-making across various domains, such as healthcare, finance, education, and law enforcement. However, concerns about fairness and bias in these systems have raised significant ethical and social challenges. To address these challenges, this research utilizes two prominent fairness libraries, Fairlearn by Microsoft and AIF360 by IBM. These libraries offer comprehensive frameworks for fairness analysis, providing tools to evaluate fairness metrics, visualize results, and implement bias mitigation algorithms. The study focuses on assessing and mitigating biases for unstructured datasets using Computer Vision (CV) and Natural Language Processing (NLP) models. The primary objective is to present a comparative analysis of the performance of mitigation algorithms from the two fairness libraries. This analysis involves applying the algorithms individually, one at a time, in one of the stages of the ML lifecycle, pre-processing, in-processing, or post-processing, as well as sequentially across more than one stage. The results reveal that some sequential applications improve the performance of mitigation algorithms by effectively reducing bias while maintaining the model’s performance. Publicly available datasets from Kaggle were chosen for this research, providing a practical context for evaluating fairness in real-world machine learning workflows.
Keywords: Machine Learning Fairness, Bias Analysis.
@@ 1 Introduction
Machine learning (ML) algorithms are extensively employed across various fields, including entertainment, shopping, healthcare, finance, education, and law enforcement, as well as in critical areas such as loan approvals [1] and hiring decisions [2, 3]. They offer benefits like consistent performance and the ability to analyze numerous variables [4, 5]. However, these systems are not immune to bias, which can lead to unfair outcomes [6, 7]. Bias in ML is particularly concerning when decisions directly affect individuals or communities, potentially resulting in discrimination. Ensuring that these systems operate ethically and equitably is crucial. Fair decision-making necessitates impartiality, avoiding favoritism based on inherent or acquired characteristics. Biased algorithms deviate from this principle, skewing decisions in favor of certain groups.
@@
The idea of ”fairness” in algorithmic systems is deeply tied to the sociotechnical context. Different forms of fairness-related harms have been identified in Table 1.
Harm Type | Definition |
Allocation Harm | Unequal distribution of opportunities or resources, such as an algorithm disproportionately favoring men over women for job offers [8]. |
Quality-of-Service Harm | Unequal performance across groups, e.g., facial recognition systems misclassifying Black women more frequently than White men [9], or speech recognition systems underperforming for users with speech disabilities [10]. |
Stereotyping Harm | Reinforcing societal stereotypes, such as image searches for ”CEO” showing predominantly White men [8]. |
Denigration Harm | Producing offensive outputs, such as misclassifying individuals as animals or chatbots generating slurs [8]. |
Representation Harm | Over- or under-representing certain groups, such as racial bias in welfare fraud investigations or neglecting elderly populations in surveillance applications [8]. |
Procedural Harm | Violations of social norms in decision-making practices, such as penalizing job candidates for extensive experience or failing to ensure transparency and accountability in algorithmic decisions [11]. |
These harms often overlap and highlight the importance of addressing fairness concerns from the development phase of ML systems.
Incorporating ML fairness techniques as an industry application in a research paper is essential for advancing ethical AI. Fairness libraries provide tools to evaluate and mitigate biases in ML models, a critical need as industries increasingly rely on AI-driven decision-making. These tools are particularly valuable for ensuring equity in sectors such as finance, banking, and healthcare. They offer intuitive and interactive ways to explore model behavior, enabling stakeholders to assess fairness trade-offs effectively. By demonstrating the practical applications of these tools, researchers can bridge the gap between academic innovation and industry adoption, fostering the development of transparent and fair AI systems. This approach not only underscores the versatility of fairness libraries but also highlights their importance in addressing real-world fairness challenges across industries.
Therefore, in this paper, we present use cases of three fairness libraries—Fairlearn (Microsoft), AIF360 (IBM), and What-If Tool (Google)—which are trusted in the industry for assessing and mitigating bias in ML models before deployment. This work aims to encourage industry professionals to implement these libraries across various applications.
In our study, we conducted a comparative evaluation of two approaches to bias mitigation in ML models. Specifically, we analyzed the effectiveness of applying a single mitigation algorithm at a time versus the sequential application of multiple mitigation algorithms across different stages of the ML lifecycle: pre-processing, in-processing, and post-processing. The sequential approach is designed to harness the strengths of stage-specific algorithms, enabling a more comprehensive bias mitigation strategy.
Figure 1 illustrates the ML lifecycle and the integration of pre-processing, in-processing, and post-processing mitigation algorithms. It depicts where a mitigation algorithm can be applied within a single ML stage (individual application) and how multiple algorithms can be implemented across different stages simultaneously (sequential application).
While many previous studies have focused on applying fairness interventions at individual stages of the ML lifecycle, our study advances the field by implementing these interventions sequentially across three stages. This comprehensive framework significantly enhances bias mitigation by addressing biases holistically, reducing the risk of fairness issues propagating through the model development process, and minimizing residual disparities that may persist with single-stage interventions [12, 13]. Applying this lifecycle-based approach to real-world datasets provides strong empirical evidence of its effectiveness. Our work bridges the gap between theoretical fairness concepts and practical implementation, encouraging broader adoption of lifecycle-based fairness practices [14]. The objectives of this study are:
@@
The central objective is to demonstrate how a sequential application of fairness mitigation algorithms across the ML lifecycle stages, pre-processing, in-processing, and post-processing, sometimes leads to good mitigation of biases compared to applying these methods in isolation.
@@
To measure the level of bias present in ML models trained on the chosen dataset.
@@
To offer actionable insights on implementing fairness tools in real-world ML workflows.
@@
This study examines the fairness of ML models applied to unstructured datasets in computer vision (CV) and natural language processing (NLP) tasks. Unstructured datasets, known for their lack of organization, facilitate efficient analysis and are commonly used in AI applications. However, these datasets often reflect historical biases or systemic discrimination, which can impact the fairness of AI models, making robust evaluation and mitigation essential. A similar fairness analysis for structured datasets was performed in our previous work [15].
@@
For this research, a publicly available Kaggle dataset [16] was selected. Kaggle datasets provide realistic scenarios for evaluating ML models and are well-suited for fairness studies. The dataset underwent preprocessing and was used to develop NLP and CV models. NLP and CV tasks are particularly significant in fairness research, as biased predictions can disproportionately affect specific groups.
@@
To identify and address biases, the study employed three advanced fairness libraries: Fairlearn by Microsoft, AIF360 by IBM, and the What-If Tool by Google. These libraries offer tools for assessing fairness metrics, visualizing model behavior, and applying bias mitigation techniques. By leveraging these tools, the research systematically evaluates the fairness of NLP and CV models and investigates strategies to minimize bias.
The remainder of this paper is structured as follows: Section 2 reviews related work on ML fairness. Section 3 details the methodology, including dataset preprocessing and model development. Section 4 discusses fairness analyses using the selected libraries and their functionalities. Section 5 presents a comparative evaluation of the libraries and their results. Future work is discussed in Section 6. Finally, Section 7 concludes with a summary of the findings.
2 Review of Related Work
Bias in ML models has become an increasingly urgent concern, particularly as these models impact critical areas such as healthcare, hiring, and criminal justice. A growing body of research has explored the origins, manifestations, and mitigation strategies for bias, providing a solid foundation for addressing this widespread challenge [17].
One key area of research focuses on identifying and categorizing biases in ML models. Reference [18] introduces a taxonomy of biases, classifying them into historical, representation, and measurement biases. Historical bias arises from pre-existing inequities in the data, even before the application of ML techniques. Representation bias occurs when certain groups are under- or overrepresented in the training data, leading to skewed predictions [19]. Measurement bias results from inaccuracies in the features or labels used for training, often caused by flawed measurement processes.
Another line of research explores methods for detecting bias. Tools such as disparate impact analysis [20] and fairness metrics like demographic parity, equal opportunity, and disparate mistreatment [21] are commonly used. In structured datasets, researchers focus on two fairness dimensions: group fairness, which ensures equitable treatment across demographic groups, and individual fairness, which treats similar individuals similarly [22]. Ref. [23] examines the conflicts between these fairness definitions, emphasizing the need for context-specific trade-offs.
The technical challenges of mitigating bias have also been extensively studied. Pre-processing techniques, such as re-weighting data samples or modifying labels to enhance fairness, address biases at the dataset level [24]. In-processing methods, such as adversarial debiasing [25], incorporate fairness constraints directly into model training. Post-processing approaches modify model outputs to meet fairness criteria, including re-ranking methods proposed by [21]. However, Ref. [26] highlights the inherent trade-offs between fairness and accuracy, demonstrating the challenge of optimizing both simultaneously.
Bias analysis in structured datasets has received significant attention due to the prevalence of tabular data in decision-making systems. These datasets often carry latent biases reflecting historical inequities or systemic discrimination. For example, the COMPAS dataset, used in criminal justice, exhibits racial disparities in predictive outcomes [27]. Research has also examined how feature selection and preprocessing can either mitigate or exacerbate biases. Ref. [28] explores the impact of feature correlation with sensitive attributes on fairness and proposes methods to disentangle these relationships.
Recent advances have emphasized interpretability as a critical component of bias analysis. Ref. [29] introduced LIME (Local Interpretable Model-agnostic Explanations), which helps stakeholders understand and detect biased patterns in model predictions. Similarly, Ref. [30] developed SHAP (SHapley Additive exPlanations), providing consistent and accurate feature importance values. These interpretability tools have been instrumental in identifying biases in structured datasets by allowing detailed analysis of how features contribute to unfair outcomes.
Multiple demographic factors have also emerged as a key consideration in bias studies. Ref. [31] emphasizes the necessity of evaluating models across various demographic dimensions, revealing compounded disparities for groups such as Black women in facial recognition systems. For structured datasets, Ref. [32] proposes fairness-enhancing strategies that account for multiple subgroups simultaneously, overcoming the limitations of single-axis fairness analyses.
The literature on bias in ML covers a wide range of topics, including foundational concepts, detection methods, mitigation strategies, and transparency tools. Despite significant progress, challenges remain in applying these techniques to structured datasets, particularly in balancing fairness with competing objectives like accuracy and transparency. This review underscores the ongoing need for research into comprehensive, context-aware methods for understanding and mitigating bias in ML models.
3 Dataset and Model Details
For the computer vision model, we used the UTKFace dataset [16], a collection of over 20,000 facial images labeled with demographic information, including age, gender, and ethnicity. Widely used for ML tasks such as facial analysis and demographic prediction, the dataset includes images cropped and aligned to ensure consistent analysis across models.
This dataset is commonly used for tasks like age and gender prediction, facial recognition, and demographic analysis. Researchers often apply deep learning techniques, including convolutional neural networks (CNNs), to extract features and train models for these predictive tasks.
For the Natural Language Processing (NLP) model, we used the California Independent Medical Review (IMR) dataset [16], available on Kaggle. This dataset contains healthcare-related data focused on medical review decisions, specifically text-based reviews related to insurance appeals. It is particularly valuable for text classification, sentiment analysis, and decision-making predictions in a healthcare context.
Although the dataset consists of text entries, researchers often use ML models such as Logistic Regression, Random Forest, or advanced NLP models like BERT to analyze decision outcomes and identify patterns in the appeals based on textual data.
For the computer vision use case, we selected the UTKFace dataset because of its broad demographic coverage. Labeled with age, gender, and ethnicity, it is especially suitable for analyzing biases in facial recognition models. Its diversity allows for a comprehensive evaluation of fairness issues, particularly regarding demographic disparities. By employing state-of-the-art facial recognition models, we aim to examine their performance across different demographic groups and explore how fairness interventions can address any observed disparities.
For the NLP use case, we chose the California Independent Medical Review dataset because it provides a rich text corpus of medical reviews, which is critical for assessing fairness in healthcare-related NLP applications. This dataset represents sensitive contexts where biased model behavior could significantly impact decision-making processes, especially in healthcare, which affects diverse patient populations. We pair this dataset with widely used NLP models to assess how fairness techniques can mitigate biases that may disproportionately impact specific groups.
We chose these datasets because they are highly relevant to real-world applications where fairness is crucial. The UTKFace dataset represents potential fairness challenges in public or commercial systems, such as those used in security or customer service. Similarly, the California Independent Medical Review dataset highlights fairness issues in healthcare, where equitable outcomes are essential. By focusing on these datasets and domains, our research aims to offer actionable insights that can be applied to similar scenarios in industry.
4 Implementation of Fairness Analyses
In recent years, the growing use of ML across various sectors has raised significant concerns regarding fairness and bias in classification models. Given their potential impact on decision-making in critical areas such as healthcare, finance, and criminal justice [33], ensuring fairness in these models is crucial. Bias in ML models can arise from factors such as imbalanced training data, model selection, and inherent societal biases, leading to discriminatory outcomes against marginalized groups [34].
In this study, we conduct a comparative evaluation to investigate the effectiveness of different strategies for mitigating bias in ML models. The methodology involves two distinct approaches:
@@ 4 Implementation of Fairness Analyses
Single-Algorithm Application: We apply individual mitigation algorithms independently at each stage of the ML lifecycle—pre-processing, in-processing, or post-processing. This approach allows us to evaluate the impact of each algorithm in isolation.
@@
Sequential Algorithm Application: We implement a sequential approach by applying multiple mitigation algorithms across the ML lifecycle stages. Specifically, we combine pre-processing, in-processing, and post-processing techniques in a systematic sequence to leverage the unique strengths of each algorithm.
For both approaches, we evaluate their performance using standard fairness metrics and monitor their impact on model accuracy. The sequential strategy aims to address bias more comprehensively by integrating mitigation efforts across different stages, ensuring a holistic reduction of bias while maintaining model performance.
This methodology provides a robust framework to compare the efficacy of standalone and sequential applications of fairness interventions, offering actionable insights for developing equitable ML systems.
@@
To address these challenges, several libraries and tools have been developed to help practitioners analyze and reduce bias in their models. Notable among them are Fairlearn, AIF360, and the What-If Tool, each offering unique features for evaluating and enhancing fairness. Details on the libraries, datasets, metrics used, and mitigation algorithms employed are provided in Table 2.
Library | Dataset | Metrics | Mitigation Algorithms |
Fairlearn by Microsoft [35] | CV: UTKFace dataset [16]. NLP: California Independent Medical Review (IMR) Dataset [16] | Demographic Parity, Equalized Odds | Correlation Remover, Exponentiated Gradient, Grid Search, and Threshold Optimizer. |
AIF360 by IBM [36] | CV: UTKFace dataset [16]. NLP: California Independent Medical Review (IMR) Dataset [16] | Statistical Parity Difference, Average Odds Difference, Equal Opportunity Difference, Theil Index, and Generalized Entropy Index. | Reweighing, Disparate Impact Remover, Learning Fair Representations, Adversarial Debiasing, Equalized Odds, and Reject Option Classification. |
What-If Tool by Google [37] | CV: UTKFace dataset [16]. NLP: California Independent Medical Review (IMR) Dataset [16] | Demographic Parity Difference, Equalized Odds Difference | N/A |
These tools are essential for researchers seeking to identify and mitigate bias in ML classification models, providing the methodologies needed to ensure equitable AI systems and informed decision-making. Table 3 summarizes the strengths and weaknesses of the libraries used in this analysis: Fairlearn and AIF360, along with the What-If Tool by Google. While the What-If Tool is valuable for model evaluation, it is limited in terms of mitigation algorithms, making it less suitable for this study.
Fairlearn | AIF360 | What-If-Tool | |
Pro | • Easy to use. Useful for educational purpose | • Large number of metrics and mitigation algorithms. • Can handle all classification tasks. | • Interactive tool with visualizations for model’s predictions. |
Cons | • Limited metrics and mitigation algorithms. • Very slow with big datasets • SUpports only binary classification tasks. | • Some algorithms needs high computational resources (e.g. Learning Fair Representation) | • Limited in terms of mitigation algorithms. • Outdated library. |